[figure: 9.7 cm]

A POSSIBLE MECHANISM OF CHARGE TRANSPORT IN COMPLEX TRANSITION METAL OXIDES

A. Lehmann-Szweykowska^{a}, R.J. Wojciechowski^{a} and G.A. Gehring^{b} aInstitute of Physics, A. Mickiewicz University ul. Umultowska 85, 61-614 Poznań, Poland bDepartment of Physics, University of Sheffield, U.K.

Group-theoretical analysis of the p-d hybridization and direct p-p hopping indicates that only a combination of both mechanisms can explain charge transport in complex transition metal oxides (eg. perovskites, garnets). The basic structural units of these crystals are MO_{4} metal-oxygen tetrahedra and/or MO_{6} octahedra. The p-d hybridized hole states are attached to each such cluster. The hole motion between the nearestneighbour clusters is mediated by the p-p symmetry-permitted hole jumps, both interand intraband. An effective hopping Hamiltonian is obtained and employed in analysis of the temperature dependence of the jump frequency. A final result turns to be a Mott-like hopping conductivity.

[figure: 13.4 cm]

Subject category :

3. Transition Metals, Alloys and Compounds

Presentation mode : oral

Corresponding author : A.Lehmann-Szweykowska

Address for correspondence : Institute of Physics,A.Mickiewicz University ul. Umultowska 85, 61-614 Poznań, Poland

Email address : als@main.amu.edu.pl